OpenStudy (anonymous):

A=1/2(bh) [solving for b]

OpenStudy (anonymous):

what are you solving for

OpenStudy (anonymous):

Solving for b

OpenStudy (anonymous):

This is the area of a triangle. Anything special? Like ... the actual question?

OpenStudy (anonymous):

Clear the fraction by multiplying both sides by the denominator which is 2. 2 * a = 2 * 1/2 * bh 2a = bh

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

there is more

OpenStudy (anonymous):

b = 2a/h

OpenStudy (anonymous):

Then you divide by 2.

OpenStudy (anonymous):

To solve for b, divide both sides by h: 2a/h = bh/h

OpenStudy (anonymous):

then your answer would be 2a/h = b

OpenStudy (anonymous):

Ohh. Okay. Thanks guys :)

OpenStudy (anonymous):

welcome